2.)Force necessary to pull a circular plate of 5cm radius from water surface for which surface tension is 75 dynes/cm, is

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
F = T * I = 75 * 2 \pi * r = 150 * 5 \pi 
 = 750 \pi dynes
Hope it helps
 
Regards
Arun (askIITians forum expert)